# Extensions of realizable Hamiltonian and complexity one GKM4 graphs

### v1

@@ 1. Introduction
GKM manifolds are a natural, far-reaching generalization of (quasi)toric manifolds, named after Goresky–Kottwitz–MacPherson [gkm-98]. These are connected, compact, orientable manifolds with vanishing odd-degree cohomology, acted on by a compact torus, in a way that the orbit space of its equivariant one-skeleton has the structure of a graph. This graph, equipped with a labelling given by the weights of the isotropy representation at the fixed points, is called the GKM graph of the action. The condition on the shape of the one-skeleton is equivalent to the weights of any isotropy representation being pairwise linearly independent; in general, if any $k$ of these weights are linearly independent, one calls the action GKMk. The starting point for this note is the following (unpublished) conjecture.
Conjecture 1 (Masuda).
@@
    Any $\operatorname{GKM}_{4}$ manifold $M^{2n}$ extends to a torus manifold. (I.e., the action extends to an effective action of $T^{n}$.)
@@
    The GKM graph $\Gamma$ of any $\operatorname{GKM}_{4}$ manifold $M^{2n}$ extends to a torus graph. (I.e., there is an abstract GKM graph with effective $T^{n}$-labeling which restricts to $\Gamma$.)
Item $(ii)$ of Conjecture 1 follows from item $(i)$; in this note we will focus on item $(ii)$. In the language of abstract GKM graphs [gu-za-01], we show (see Theorem 4.19):
@@
Our main tool for the proof of Theorem 2 is Kuroki’s axial function group [ku-19] and his criterion for the existence of nontrivial extensions of labelings of GKM graphs. In Section 4 we extend Kuroki’s results to unsigned GKM graphs, and use index-free notation to rewrite the axial function group as the group of invariants under a natural representation of the fundamental group $\pi_{1}(\Gamma)$.
@@
Combining this statement with Theorem 2, it follows that the GKM graph of any Hamiltonian GKM4 manifold extends to a torus graph. This result proves a particular case of Conjecture 1 $(ii)$, under the additional Hamiltonian assumption. Theorem 3 follows from a more general statement about GKM3 graphs with an acyclic orientation of the edges, so that every $2$-face has a unique local maximum with respect to the orientation, see Theorem 3.9.
The second situation concerns GKM4 manifolds of complexity $1$, meaning that we are given an effective action of an torus of dimension n1n-1 on a manifold of dimension $2n$.
@@
Theorem 4 proves another particular case for Conjecture 1 $(ii)$, namely, in complexity $1$. To prove Theorem 4 (see Theorem 5.6) we construct a certain GKM cover $p\colon\widetilde{\Gamma}\to\Gamma$ by a (possibly infinite) GKM graph $\widetilde{\Gamma}$ whose conjugated $2$-faces generate the respective fundamental group. Then we apply Theorem 2 to obtain an extension of $\widetilde{\Gamma}$ to a torus graph. We conclude by showing that the action of the deck transformation group on the axial function group of $\widetilde{\Gamma}$ is trivial, making use of the fact that it is a perfect group by the acyclicity results for orbit spaces of equivariant skeleta in [ay-ma-23]. Notice that Theorem 4 does not hold for abstract GKM graphs by [so-23].
Acknowledgements: The authors gratefully acknowledge funding of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation): Project number 452427095, as well as the second named author’s Walter Benjamin Fellowship. The second named author is grateful to S. Kuroki and M. Masuda (in particular for organizing the conference “Toric Topology in Okayama 2019”, where he was happy to participate and present some of the ideas from the current note for the first time). We wish to thank P. Konstantis, L. Zoller and N. Wardenski for various stimulating discussions on the subject of this note.
@@ 2. Graphs, connections, invariant functions
We will need to keep track of the signs $\varepsilon$ and the integer coefficients in the above congruence relation. To this end we fix, for easier bookkeeping, an arbitrary global lift
@@
of $\alpha$ (i.e., $\pm\widetilde{\alpha}(e)=\alpha(e)$ for all $e\in E$), satisfying no further conditions except $\widetilde{\alpha}(\overline{e})=-\widetilde{\alpha}(e)$ for all $e\in E$.
In order to avoid having to choose an ordering of the edges at each vertex, we introduce the following notation. Define a nondegenerate scalar product on $\mathbb{Z}\operatorname{star}v$ by requiring $\operatorname{star}v$ to be an orthonormal basis, i.e.,
@@
Denote the group $\{\pm 1\}$ by $\mu_{2}$. Let $\mu_{2}\operatorname{star}v:=(\mu_{2})^{\operatorname{star}v}$ be the group of all $\mu_{2}$-valued functions on the finite set $\operatorname{star}v$ with respect to pointwise group structure. This group acts on $\mathbb{Z}\operatorname{star}v$ by the formula below:
$\mu_{2}\operatorname{star}v\times\mathbb{Z}\operatorname{star}v\to\mathbb{Z}\operatorname{star}v,\ x\cdot y:=\sum_{e^{\prime}\in\operatorname{star}v}\langle x,e^{\prime}\rangle\langle y,e^{\prime}\rangle e^{\prime},$
which is nothing but componentwise multiplication in the standard basis.
@@
    The signed axial function $\alpha$ is compatible with the conection, i.e., for all edges $e,e^{\prime}$ at any vertex $v\in V$,
@@ 3.1. Definitions and basic facts
In what follows by a face we mean either a face submanifold, face of an orbit space or a face of a GKM graph, whenever it is clear from the context. By the following lemma, the condition on the existence of a fixed point (i.e. of a vertex of a face) in Definition 3.1 is superfluous for GKM manifolds. (The claim of this lemma is is a particular case of [ma-pa-06, Lemma 2.2] in the case of GKM manifolds, specifically, having only finitely many $T$-fixed points.)
@@ 3.2. Proof of Theorem 3
Given any cycle $g$ at the least vertex $v_{0}$ of $\Gamma$, we wish to decompose $g$ into a product of conjugated $2$-faces. This is done by double induction on height and multiplicity of $g$. Namely, let $v$ be the maximal vertex of $g$, and consider one of its occurrences, with adjacent edges $uv$, vwvw.
@@
Notice that $g_{1}$ is a conjugated $2$-face corresponding to $F$, and $g_{2}$ is a path with lower multiplicity or height, as it traverses $F$ from the other side avoiding $v$, see Fig. 1. We can therefore apply the induction hypothesis to $g_{2}$.
@@
Figure 1. The induction step replaces the edge path $u,v,w$ with the (dashed) edge path $u,z,w$ in a cycle $g$ with a maximal vertex $v$ of multiplicity 22. The vertices $u,v,w,z$ belong to the same 22-face $F$. The resulting cycle has the maximal vertex $v$ with multiplicity 11.
@@
Corollary 3.11 is not true in the GKM2 case, and Theorem 3.9 does not hold without the assumption on the Betti numbers of the $2$-faces. For example, the standard $T^{2}$-action on the flag manifold $Fl_{3}={\mathrm{SU}}(3)/T^{2}$ has the GKM graph $\Gamma$ as in Figure 2. As the action is Hamiltonian, a generic component of the momentum map induces an acyclic orientation on $\Gamma$. Equipped with the canonical connection of a homogeneous GKM manifold [gu-ho-za-06], it has a $2$-face $F$ consisting of a $6$-cycle subgraph that satisfies $b_{4}(F)=2$. On the other hand, the minimal normal subgroup generated by conjugated $2$-faces of $\Gamma$ has index $2$ in $\pi_{1}(\Gamma)$. This follows for instance because gluing in a $2$-disc to each $2$-face of $\Gamma$ results in $\mathbb{R}P^{2}$, see [go-ko-zo-22', Example 3.13 (b)]
Figure 2. GKM graph of the flag manifold $Fl_{3}={\mathrm{SU}}(3)/T^{2}$. It has a $2$-face $F$ with $b_{4}(F)=2$, depicted by dotted lines.
@@ 4.2. Group of axial functions
$=\sum_{e^{\prime}\in\operatorname{star}i(e)}\langle\varepsilon(e),e^{\prime}\rangle\cdot a(e^{\prime})\cdot\nabla_{e}e^{\prime}+\langle c(e),e^{\prime}\rangle\cdot a(e)\cdot\nabla_{e}e^{\prime}$
$=\sum_{e^{\prime}\in\operatorname{star}i(e)}\langle\varepsilon(e),e^{\prime}\rangle\cdot a(e^{\prime})\cdot\nabla_{e}e^{\prime}+a(e)\cdot\nabla_{e}c(e)$
@@
(7) $\varphi_{e}(e)=\overline{e}+\nabla_{e}c(e)=\overline{e}+\varepsilon(\overline{e})\cdot c(\overline{e}),\qquad\varphi_{e}(e^{\prime})=\nabla_{e}e^{\prime}\quad\textrm{for}\quad e^{\prime}\neq e.$
@@
the A-monodromy representation. The image of $\pi_{1}(\Gamma,u)$ in $\operatorname{GL}(\mathbb{Z}\operatorname{star}u)$ is called the A-monodromy group at $u\in\Gamma$.
@@
The A-monodromy representation is well-defined.
@@
$\varphi_{\overline{e}}\circ\varphi_{e}(e^{\prime})=\varphi_{\overline{e}}(\nabla_{e}e^{\prime})=\nabla_{\overline{e}}\nabla_{e}e^{\prime}=e^{\prime}.$
@@ 4.3. Extension criterion for unsigned GKM graphs
Figure 3. The signed GKM graph (with the connection acting by nontrivial permutation along every edge) of the canonical $T^{2}$-action on $G_{2}/SU_{3}=S^{6}$ does not extend to a torus graph [ku-19]; the respective unsigned GKM graph (with the different connection sending every edge to itself) extends to an unsigned torus graph, corresponding to the standard action of $T^{3}$ on $S^{6}\subset\mathbb{C}^{3}\oplus\mathbb{R}$.
@@ 4.4. Proof of Theorem 2
Since the original labeling of $\Gamma$ induces on $\Gamma^{\prime}$ the structure of a torus graph, $A(\Gamma)\cong(\mathbb{Z}\operatorname{star}v)^{\pi_{1}(\Gamma,v)}$ contains two linearly independent elements that remain linearly independent upon projection to
$\mathbb{Z}\operatorname{star}_{\Gamma^{\prime}}v=\mathbb{Z}e_{1}\oplus\mathbb{Z}\overline{e_{q}},$
where $\gamma$ follows the edge path $(e_{1},\ldots,e_{q})$. In particular, $\varphi_{\gamma}$ acts trivially on these two elements.
@@ 5. Extensions of realizable complexity one GKM4 graphs
$\sum_{i=1}^{n}a_{i}(gh)f_{i}=g\cdot(h\cdot f_{n})=\sum_{i=1}^{n-1}a_{i}(h)g\cdot f_{i}+g\cdot f_{n}=\sum_{i=1}^{n}a_{i}(h)f_{i}+\sum_{i=1}^{n}a_{i}(g)f_{i}.$
@@
Consider a more general situation (related to Conjecture 1 $(ii)$), when $\Gamma$ is any (finite) abstract $\operatorname{GKM}_{4}$ graph. Then there exists a GKM cover $\pi\colon\widetilde{\Gamma}\to\Gamma$ that is trivial over $2$-faces such that its fundamental group is generated by conjugated $2$-faces. Then $\widetilde{\Gamma}$ extends to a torus graph by Theorem 4.19. To prove this claim glue in $2$-discs to every $2$-face in $\Gamma$ and consider the universal covering of the resulting space as described above, as well as $\widetilde{\Gamma}:=\pi^{-1}(\Gamma)$. In this situation, however, we cannot apply Corollary 3.6.

### v2

@@ 1. Introduction
GKM manifolds are a natural, far-reaching generalization of (quasi)toric manifolds, named after Goresky–Kottwitz–MacPherson [gkm-98]. These are connected, compact, orientable manifolds with vanishing odd-degree cohomology, acted on by a compact torus, in a way that the orbit space of its equivariant one-skeleton has the structure of a graph. This graph, equipped with a labelling given by the weights of the isotropy representation at the fixed points, is called the GKM graph of the action. The condition on the shape of the one-skeleton is equivalent to the weights of any isotropy representation being pairwise linearly independent; in general, if any $k$ of these weights are linearly independent, one calls the action GKMk. The starting point for this note is the following problem.
Problem 1 (Masuda).
@@
    Does any $\operatorname{GKM}_{4}$ manifold $M^{2n}$ extend to a torus manifold? (I.e., does the action extend to an effective action of $T^{n}$?)
@@
    Does the GKM graph $\Gamma$ of any $\operatorname{GKM}_{4}$ manifold $M^{2n}$ extend to a torus graph? (I.e., is there an abstract GKM graph with effective $T^{n}$-labeling which restricts to $\Gamma$?)
A positive answer to item $(i)$ of Problem 1 implies a positive answer to item $(ii)$; in this note we will focus on item $(ii)$. In the language of abstract GKM graphs [gu-za-01], we show (see Theorem 4.19):
@@
Our main tool for the proof of Theorem 2 is Kuroki’s axial function group [ku-19] and his criterion for the existence of nontrivial extensions of labelings of GKM graphs. The group of axial functions was defined in [ku-19] for signed GKM graphs (e.g., those of GKM manifolds with an invariant almost complex structure). A signed GKM graph yields an unsigned GKM graph by composing the axial function with the quotient of the lattice by $\{\pm 1\}$. In Section 4 we extend Kuroki’s results to unsigned GKM graphs, and use index-free notation to rewrite the axial function group as the group of invariants under a natural representation of the fundamental group $\pi_{1}(\Gamma)$.
@@
Combining this statement with Theorem 2, it follows that the GKM graph of any Hamiltonian GKM4 manifold extends to a torus graph. This result positively resolves a particular case of Problem 1 $(ii)$, under the additional Hamiltonian assumption. Theorem 3 follows from a more general statement about GKM3 graphs with an acyclic orientation of the edges, so that every $2$-face has a unique local maximum with respect to the orientation, see Theorem 3.10.
The second situation concerns GKM4 manifolds of complexity $1$, meaning that we are given an effective action of a torus of dimension n1n-1 on a manifold of dimension $2n$.
@@
Theorem 4 positively resolves another particular case for Problem 1 $(ii)$, namely, in complexity $1$. To prove Theorem 4 (see Theorem 5.6) we construct a certain GKM cover $p\colon\widetilde{\Gamma}\to\Gamma$ by a (possibly infinite) GKM graph $\widetilde{\Gamma}$ whose conjugated $2$-faces generate the respective fundamental group. Then we apply Theorem 2 to obtain an extension of $\widetilde{\Gamma}$ to a torus graph. We conclude by showing that the action of the deck transformation group on the axial function group of $\widetilde{\Gamma}$ is trivial, making use of the fact that it is a perfect group by the acyclicity results for orbit spaces of equivariant skeleta in [ay-ma-23]. Notice that Theorem 4 does not hold for abstract GKM graphs by [so-23].
Acknowledgements: The authors gratefully acknowledge funding of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation): Project numbers 452427095, and 561158824 (the second named author’s Walter Benjamin Fellowship). The second named author is grateful to S. Kuroki and M. Masuda (in particular for organizing the conference “Toric Topology in Okayama 2019”, where he was happy to participate and present some of the ideas from the current note for the first time). We wish to thank P. Konstantis, L. Zoller and N. Wardenski for various stimulating discussions on the subject of this note.
@@ 2. Graphs, connections, invariant functions
We will need to keep track of the signs $\varepsilon$ and the integer coefficients in the above congruence relations. To this end we fix, for easier bookkeeping, an arbitrary global lift
@@
of $\alpha$ (i.e., $\pi\circ\widetilde{\alpha}=\alpha$, where $\pi:\mathbb{Z}^{k}\to\mathbb{Z}^{k}/\pm 1$ is the canonical projection), satisfying no further conditions except $\widetilde{\alpha}(\overline{e})=-\widetilde{\alpha}(e)$ for all $e\in E$. This lift $\widetilde{\alpha}$ is not to be confused with the choice of a signed structure on the GKM graph (see Definition 2.7 below).
In order to avoid having to choose an ordering of the edges at each vertex, we introduce the following notation. Define a nondegenerate scalar product on the free $\mathbb{Z}$-module $\mathbb{Z}\operatorname{star}v$ generated by the finite set $\operatorname{star}v$ by requiring $\operatorname{star}v$ to be an orthonormal basis, i.e.,
@@
Denote the group $\{\pm 1\}$ by $\mu_{2}$. Let $\mu_{2}\operatorname{star}v$ be the free $\mu_{2}$-module generated by the finite set $\operatorname{star}v$. This group acts on $\mathbb{Z}\operatorname{star}v$ by the formula below:
$\mu_{2}\operatorname{star}v\times\mathbb{Z}\operatorname{star}v\to\mathbb{Z}\operatorname{star}v,\ x\cdot y:=\sum_{e\in\operatorname{star}v}\langle x,e\rangle\langle y,e\rangle e,$
which is nothing but componentwise multiplication in the standard basis. Here, we wrote $x\in\mu_{2}\operatorname{star}v$ in the standard basis as $x=\sum_{e\in\operatorname{star}v}\langle x,e\rangle e$, with $\langle x,e\rangle\in\mu_{2}$. In a similar way to the above case of $\mathbb{Z}$-modules, we extend $\nabla_{e}$ uniquely to a homomorphism of free $\mu_{2}$-modules $\mu_{2}\operatorname{star}i(e)\to\mu_{2}\operatorname{star}t(e)$.
@@
    The signed axial function $\alpha$ is compatible with the connection, i.e., for all edges $e,e^{\prime}$ at any vertex $v\in V$,
@@ 3.1. Definitions and basic facts
In what follows by a face we mean either a face submanifold, face of an orbit space or a face of a GKM graph, whenever it is clear from the context. By the following lemma, the condition on the existence of a fixed point (i.e. of a vertex of a face) in Definition 3.1 is superfluous for GKM manifolds. (The claim of this lemma is a particular case of [ma-pa-06, Lemma 2.2] in the case of GKM manifolds, specifically, having only finitely many $T$-fixed points.)
@@ 3.2. Proof of Theorem 3
Remark 3.9.
Suppose $\Gamma$ is a signed GKM graph, with axial function $\alpha\colon E\to\mathbb{Z}^{k}$. For a generic vector $\xi\in(\mathbb{Z}^{k})^{*}$ the pairing $\langle\alpha(e),\xi\rangle$ is nonzero and induces an orientation $o=o_{\xi}$ [gu-za-01, §1.4]. The respective numbers $b_{2i}(\Gamma,o_{\xi})$ are called combinatorial Betti numbers and denoted by $b_{2i}(\Gamma)$ in [gu-za-01, §1.4]. As the notation suggests, $b_{2i}(\Gamma)$ does not depend on the choice of $\xi\in(\mathbb{Z}^{k})^{*}$. Furthermore, in the Hamiltonian setting $\Gamma=\Gamma(M)$, one has $b_{2i}(\Gamma)=b_{2i}(M)$ by [gu-za-01, Formula (0.4)], which explains the nomenclature. For an arbitrary (acyclic) orientation $o$ on edges of $\Gamma$, $b_{2i}(\Gamma,o_{\xi})$ depends on the choice of $o$, in general (see Fig. 1), and such numbers do not relate to topological Betti numbers.
Figure 1. The same GKM graph with two different acyclic edge orientations providing different fourth Betti numbers (i.e. numbers of vertices that are local maxima with respect to a given orientation)
@@
Given any cycle $g$ at the least vertex $v_{0}$ of $\Gamma$, we wish to decompose $g$ into a product of conjugated $2$-faces. This is done by double induction on height and multiplicity of $g$. The double induction is done with respect to the lexicographical order $\leqslant_{lex}$ on $(h,m)\in\mathbb{N}\times\mathbb{N}$. The base of the induction is $(h,m)=(1,1)$. In this case, the edge path $g$ is constant and there is nothing to prove. Assume that the claim holds for $\leqslant_{lex}(h-1,m)$ and $\leqslant_{lex}(h,m-1)$. Let $g$ be an edge path with height and multiplicity $(h,m)$. Let $v$ be the maximal vertex of $g$, and consider one of its occurrences, with adjacent edges $uv$, vwvw.
@@
Notice that $g_{1}$ is a conjugated $2$-face corresponding to $F$, and $g_{2}$ is a path with lower multiplicity or height, as it traverses $F$ from the other side avoiding $v$, see Fig. 2. Therefore, $g_{2}$ decomposes by the induction assumption. Hence, $g$ decomposes as well. This completes the induction step, and the proof is complete.
@@
Figure 2. The induction step replaces the edge path $u,v,w$ with the (dashed) edge path $u,z,w$ in a cycle $g$ with a maximal vertex $v$ of multiplicity $2$. The vertices $u,v,w,z$ belong to the same 22-face $F$. The resulting cycle has the maximal vertex $v$ with multiplicity 11
@@
Corollary 3.12 is not true in the GKM2 case, and Theorem 3.10 does not hold without the assumption on the Betti numbers of the $2$-faces. For example, the standard $T^{2}$-action on the flag manifold $Fl_{3}={\mathrm{SU}}(3)/T^{2}$ has the GKM graph $\Gamma$ as in Figure 3. As the action is Hamiltonian, a generic component of the momentum map induces an acyclic orientation on $\Gamma$. Equipped with the canonical connection of a homogeneous GKM manifold [gu-ho-za-06], it has a $2$-face $F$ consisting of a $6$-cycle subgraph that satisfies $b_{4}(F)=2$. On the other hand, the minimal normal subgroup generated by conjugated $2$-faces of $\Gamma$ has index $2$ in $\pi_{1}(\Gamma)$. This follows for instance because gluing in a $2$-disc to each $2$-face of $\Gamma$ results in $\mathbb{R}P^{2}$, see [go-ko-zo-22', Example 3.13 (b)].
Figure 3. GKM graph of the flag manifold $Fl_{3}={\mathrm{SU}}(3)/T^{2}$. It has a $2$-face $F$ with $b_{4}(F)=2$, depicted by dotted lines
@@ 4.2. Group of axial functions
$=\sum_{e^{\prime}\in\operatorname{star}i(e)}(\langle\varepsilon(e),e^{\prime}\rangle\cdot a(e^{\prime})\cdot\nabla_{e}e^{\prime}+\langle c(e),e^{\prime}\rangle\cdot a(e)\cdot\nabla_{e}e^{\prime})$
$=\biggl(\sum_{e^{\prime}\in\operatorname{star}i(e)}\langle\varepsilon(e),e^{\prime}\rangle\cdot a(e^{\prime})\cdot\nabla_{e}e^{\prime}\biggr)+a(e)\cdot\nabla_{e}c(e)$
@@
(7) $\varphi_{e}(e)=\overline{e}+\nabla_{e}c(e)=\overline{e}+\varepsilon(\overline{e})\cdot c(\overline{e}),\qquad\varphi_{e}(e^{\prime})=\varepsilon(\overline{e})\cdot\nabla_{e}e^{\prime}\quad\textrm{for}\quad e^{\prime}\neq e.$
@@
the A-monodromy representation, where $\gamma]$ denotes the class in $\pi_{1}(\Gamma,u)$ of an edge loop $\gamma$ based at the vertex $u$ in $\Gamma$. The image of $\pi_{1}(\Gamma,u)$ in $\operatorname{GL}(\mathbb{Z}\operatorname{star}u)$ is called the A-monodromy group at $u\in\Gamma$.
@@
The A-monodromy representation is well-defined. More precisely: Let $\gamma$, $\gamma^{\prime}$ be any two edge loops based at $u$ in $\Gamma$ such that $\gamma]=[\gamma^{\prime}]$ holds in $\pi_{1}(\Gamma,u)$. Then one has $\varphi_{\gamma}=\varphi_{\gamma^{\prime}}$.
@@
$\varphi_{\overline{e}}\circ\varphi_{e}(e^{\prime})=\varphi_{\overline{e}}(\varepsilon(\overline{e})\cdot\nabla_{e}e^{\prime})=\varepsilon(\overline{e})\cdot\varepsilon(e)\cdot\nabla_{\overline{e}}\nabla_{e}e^{\prime}=e^{\prime}.$
@@ 4.3. Extension criterion for unsigned GKM graphs
Figure 4. The signed GKM graph (with the connection acting by nontrivial permutation along every edge) of the canonical $T^{2}$-action on $G_{2}/SU_{3}=S^{6}$ does not extend to a torus graph [ku-19]; the respective unsigned GKM graph (with the different connection sending every edge to itself) extends to an unsigned torus graph, corresponding to the standard action of $T^{3}$ on $S^{6}\subset\mathbb{C}^{3}\oplus\mathbb{R}$
@@ 4.4. Proof of Theorem 2
Since the original labeling of $\Gamma$ induces on $\Gamma^{\prime}$ the structure of a torus graph, $A(\Gamma)\cong(\mathbb{Z}\operatorname{star}v)^{\pi_{1}(\Gamma,v)}$ contains two linearly independent elements, say $x_{i}$, $i=1,2$. By definition of the action of $\pi_{1}(\Gamma,v)$, these are in particular invariant under $\varphi_{\gamma}$. Furthermore, $x_{i}$, $i=1,2$, remain linearly independent upon the natural projection
$\pi\colon\mathbb{Z}\operatorname{star}_{\Gamma}v\to\mathbb{Z}\operatorname{star}_{\Gamma^{\prime}}v=\mathbb{Z}e_{1}\oplus\mathbb{Z}\overline{e_{q}}.$
@@ 5. Extensions of realizable complexity one GKM4 graphs
$\sum_{i=1}^{n}a_{i}(gh)f_{i}=g\cdot(h\cdot f_{n})=\sum_{i=1}^{n-1}a_{i}(h)g\cdot f_{i}+g\cdot f_{n}=\sum_{i=1}^{n-1}a_{i}(h)f_{i}+\sum_{i=1}^{n}a_{i}(g)f_{i}.$
@@
Consider a more general situation (related to Problem 1 $(ii)$), when $\Gamma$ is any (finite) abstract $\operatorname{GKM}_{4}$ graph. Then there exists a GKM cover $\pi\colon\widetilde{\Gamma}\to\Gamma$ that is trivial over $2$-faces such that its fundamental group is generated by conjugated $2$-faces. Then $\widetilde{\Gamma}$ extends to a torus graph by Theorem 4.19. To prove this claim glue in $2$-discs to every $2$-face in $\Gamma$ and consider the universal covering of the resulting space as described above, as well as $\widetilde{\Gamma}:=\pi^{-1}(\Gamma)$. In this situation, however, we cannot apply Corollary 3.6.